# When Is Diversity Rewarded in Cooperative Multi-Agent Learning?

### v1

@@ 3 Analysis
4 Heterogeneous Environment Design (HED)
In complex scenarios where theory might be less applicable, we study heterogeneity through algorithmic search. We consider the setting of a Parametrized Dec-POMDP (PDec-POMDP, defined in App. K). A PDec-POMDP represents a Dec-POMDP oliehoek2016concise , where the observations, transitions, or reward are conditioned on parameters $\theta$. Hence, the return obtained by the agents, $G^{\theta}(\mathbf{\pi})$, can be differentiated with respect to $\theta$: $\nabla_{\theta}G^{\theta}(\mathbf{\pi})=\frac{\partial}{\partial_{\theta}}G^{\theta}(\mathbf{\pi})$. In particular, computing this gradient in a differentiable simulator allows us to back-propagate through time and optimize $\theta$ via gradient ascent [^3] Although the same approach can train policies xu2022accelerated ; song2024learning , HED instead optimizes environment parameters and policies separately, using standard zeroth-order policy-gradient methods, to avoid being trapped in local minima..
Algorithm 1 Heterogeneous Environment Design (HED).
@@ 4 Heterogeneous Environment Design (HED)
[…]
5:     if train_env($i$then
@@
7:     if train_agents($i$then
@@
We now consider the problem of learning the environment parameters $\theta$ to maximize the empirical heterogeneity gain. The empirical heterogeneity gain is defined as the difference in performance between heterogeneous and homogeneous teams in a given PDec-POMDP parametrization. Heterogeneous agents each learn different parameters, whereas homogeneous agents learn the same policy. We denote their policies as $\mathbf{\pi}_{\mathrm{het}}$ and $\mathbf{\pi}_{\mathrm{hom}}$. Then, we can simply write the gain as: $\mathrm{HetGain}^{\theta}(\mathbf{\pi}_{\mathrm{het}},\mathbf{\pi}_{\mathrm{hom}})=G^{\theta}(\mathbf{\pi}_{\mathrm{het}})-G^{\theta}(\mathbf{\pi}_{\mathrm{hom}})$, representing the return of heterogeneous agents minus that of homogeneous agents on environment parametrization $\theta$. HED, shown in Alg. 1, learns $\theta$ by performing gradient ascent to maximize the gain: $\theta\leftarrow\theta+\alpha\nabla_{\theta}\mathrm{HetGain}^{\theta}(\mathbf{\pi}_{\mathrm{het}},\mathbf{\pi}_{\mathrm{hom}})$. The environment and the agents are trained in an iterative process. At every training iteration, HED collects roll-out batches in the current environment $\theta$ for both heterogeneous and homogeneous teams, computing the heterogeneity gain on the collected data. Then, it updates $\theta$ to maximize the heterogeneity gain. Finally, to train the agents, it uses MARL, with any on-policy algorithm (e.g., MAPPO yu2022surprising ). The functions train_env and train_agents determine when to train each of the components in HED. In particular, we consider two possible training regimes: (1) alternated: where HED performs cycles of $x$ agent training iterations followed by $y$ environment training iterations and (2) concurrent: where agents train at every iteration and the environment is updated every $x$ iterations. Note that by performing descent instead of ascent, HED can also be used to minimize the heterogeneity gain, which may be useful in homogeneous systems.
@@ 5 Experiments
To empirically ground our theoretical analysis, we conduct a three-stage experimental study in cooperative MARL. We first analyze a one-step, observation-free matrix game in which each agent allocates effort $r_{ij}$ over $M$ tasks, and consider reward structures defined by aggregator pairs $U,T\in\{\min,\mathrm{mean},\max\}$. We find that the agents’ learned policies recover the exact heterogeneity gains derived in the theory (Fig. 2). Next, we transfer the same reward structures into an embodied, time-extended Multi-Goal-Capture environment. We show that our curvature theory continues to be predictive in this setting as well, and discuss interesting learning dynamics emerging from the agents’ embodiment. We also expose a trade-off between the heterogeneity gain and agents’ observations, showing that when homogeneous agents can access richer observations, they can learn incrementally heterogeneous behaviors by leveraging behavioral typing and thus minimizing the gain: this highlights the difference between neural and behavioral heterogeneity bettini2023hetgppo .
Finally, to study HED, we parametrize the reward structure of Multi-Goal-Capture using either parametrized Softmax or Power-Sum aggregators (App. H), and run HED to learn parameterizations that maximize the heterogeneity gain. HED learns the theoretically optimal aggregator instantiations: one the one hand, this validates HED’s effectiveness at discovering heterogeneous missions, and on the other hand, showcases the generalizability of our theoretical insights to embodied environments. Implementation details and visualizations are available in App. C and App. J, respectively.
@@
We consider a one-step observationless matrix game where $N$ agents need to choose between $M$ tasks. Their actions are effort allocations $r_{ij}$ with $r_{ij}\geq 0$,$\sum_{j}r_{ij}=1$, composing matrix $\mathbf{A}$. With aggregators taken from the set $U,T\in\{\min,\mathrm{mean},\max\}$, our goal is to empirically confirm the heterogeneity gains derived in the theory in a learning context. Each time the game is played, all agents obtain the global reward $R(\mathbf{A})$ computed through the double aggregator. We consider two setups: (1) Continuous ($r_{ij}\in\mathbb{R}_{0\leqslant x\leqslant 1}$): agents can distribute their efforts across tasks, (2) Discrete ($r_{ij}\in\{0,1\}$): agents choose only one task. We train with $N=M=4$ for 12 million steps. The evolution of the heterogeneity gains over training is shown in Fig. 6. The final results match exactly the theoretical predictions in the Fig. 2 table and our curvature theory: concave outer and convex inner aggregators favor heterogeneity. Additional results and details are in App. I.
@@
Figure 3: Heterogeneity gain for the discrete and continuous matrix games with $N=M=4$ over training iterations. We report mean and standard deviation after 12M frames over 9 random seeds. The final results match the theoretical predictions in the Table of Fig. 2.
Next, we investigate a time-extended, embodied scenario called Multi-Goal-Capture, based on multi-goal navigation missions terry2021pettingzoo . Our goal is to show that the results obtained in the matrix game still hold in this more complex setting and that aggregators that maximize the heterogeneity gain encode a practically useful global objective.
In Multi-Goal-Capture, agents need to navigate to goals (visualization availabile in App. J). Each agent observes the relative position to the goals, and agent actions are continuous 2D forces that determine their direction of motion. The entries $r^{t}_{ij}$ of matrix $\mathbf{A}^{t}$ at time $t$ represent the local reward of agent $i$ towards goal $j$, computed as $r^{t}_{ij}=\left(1-d^{t}_{ij}/\sum_{j=1}^{M}d^{t}_{ij}\right)/(M-1),$ where $d^{t}_{ij}$ is the distance between agent $i$ and goal $j$. This makes it so that $\sum_{j=1}^{M}r^{t}_{ij}=1$ and $r^{t}_{ij}\geqslant 0$. At each step, the agents receive the global reward $R(\mathbf{A}^{t})$, again with aggregators $U,T\in\{\min,\mathrm{mean},\max\}$. Different aggregator choices will yield different global objectives. For example, $U=\max,T=\max$ implies “at least one agent should go to at least one goal”; $U=\max,T=\min$ implies “all agents should go to the same goal”, and so on. $U=\min,T=\max$, a concave-convex setting shown by our theory to favor heterogeneity, implies “each agent should go to a different goal and all goals should be covered” which is a natural goal for this scenario. This is because $T=\max$ encodes a task that needs just one agent to be completed (e.g., find an object), while $U=\min$ encodes that all tasks should be attended (i.e., agents need to diversify their choices).
(a) Gain for different aggregators.
(b) Gain w.r.t. observability when $U=\min,T=\max$.
@@
Figure 4: Heterogeneity gain for Multi-Goal-Capture throughout training. We report mean and standard deviation for 30 million training frames over 9 (left), 4 (right) random seeds.
After 30M environment frames (Fig. 4(a)) the empirical heterogeneity gains differ, numerically, from those of the static matrix-game because agents now realize their allocations $r_{ij}$ through time-extended motion. Nonetheless, our curvature theory still predicts when there is a heterogeneity gain (Fig. 2): it is positive only for the concave–convex pairs $U=\min,\,T=\max$ and $U=\mathrm{mean},\,T=\max$. The heterogeneity gain is smaller in the latter case because learning dynamics matter: with $U=\min,\,T=\max$ the best homogeneous policy is unique–every agent must steer to the midpoint between the two goals–so homogeneous learners seldom find it, leaving room for heterogeneous policies to excel (see App. J). By contrast, $U=\mathrm{mean},\,T=\max$ admits a continuum of good homogeneous policies, which homogeneous teams execute more easily. For $U=\max,\,T=\min$ and $U=\max,\,T=\mathrm{mean}$ the empirical heterogeneity gap is actually negative: this is because the reward peaks only when all agents choose the same goal, a coordination that heterogeneous teams learn more slowly, so they lag within the fixed training budget. Crucially, while additional training will eventually result in $\Delta R=0$ in the latter case, the positive heterogeneity gains we report are theoretically irreducible, pinpointing scenarios where behavioral diversity is needed.
Observability-Heterogeneity Trade-Off: We now crystallize the relationship between environment observability and empirical heterogeneity gains. It is well known that neurally homogeneous agents (i.e., sharing the same neural network) can emulate diverse behavior by conditioning on the input context (behavioral typing) leibo2019malthusian . This can be achieved by naively appending the agent index to its observation gupta2017cooperative or by providing relevant observations that allows the agents to infer their role bettini2023hetgppo . Behavioral typing is impossible in matrix games, as these games are observationless. However, it is possible in more complex games, such as our Multi-Goal-Capture scenario. We augment agents in the positive gain scenario ($U=\min,T=\max$) with a range sensor, providing proximity readings for other agents within a radius. In 4(b), we show that the heterogeneity gain decreases as the agent visibility increases (higher sensing radius). This is because, with a higher range, homogeneous agents can sense each other and coordinate to pursue different goals. This result highlights the tight interdependence between the heterogeneity gain and agents’ observations.
@@
We apply HED to Multi-Goal-Capture, and ask whether it finds the same aggregator parameterizations predicted by our theory to maximize the heterogeneity gain. We turn the environment into a PDec-POMDP by parameterizing the reward as ${R}^{\theta}(\mathbf{A}^{t})=\bigoplus_{j=1}^{M}\!{}^{\theta}\bigoplus_{i=1}^{N}\!{}^{\theta}r^{t}_{ij}$, with parametrized inner and outer aggregators $U^{\theta}=\bigoplus^{\theta},T^{\theta}=\bigoplus^{\theta}$. Our goal is to learn the parameters $\theta=(\tau_{1},\tau_{2})$, parametrizing $T$ and $U$ respectively, that maximize the heterogeneity gain. We consider two parametrized aggregators from Tab. 3: Softmax and Power-Sum. Softmax: We parameterize both $U^{\theta}$ and $T^{\theta}$ using Softmax. We initialize $\tau_{1}=\tau_{2}=0$, so $U$ and $T$ are initially mean, and run HED. In 5(a), we show that, to maximize the heterogeneity gain, HED learns to maximize $\tau_{1}$, making the inner aggregator $T$ Schur-convex, while minimizing $\tau_{2}$, making the outer aggregator $U$ Schur-concave. The result is exactly the reward function predicted by our theory to maximize the gain. Power-Sum: We parametrize both aggregators with Power-Sum. We initialize both functions to $\tau_{1}=\tau_{2}=1$, representing sum, and run HED. We constrain $\tau_{1,2}\in[0.3,6]$ to stabilize learning. In 5(b), we show that HED learns to maximize $\tau_{1}$, making $T$ Schur-convex, while minimizing $\tau_{2}$, making $U$ Schur-concave; again rediscovering the optimal parametrization our theory predicts. Hence, these results simultaneously validate HED and our curvature theory.
@@
(a) HED with Softmax aggregators ($\tau\in\mathbb{R}$). $\tau\geq 0$ is Schur-convex; $\tau\leq 0$ is Schur-concave.
(b) HED with Power-Sum aggregators ($\tau\in[0.3,6]$). $\tau\geq 1$ is Schur-convex; $\tau\leq 1$ is Schur-concave.
@@
Figure 5: HED results in Multi-Goal-Capture. The two leftmost columns report the evolution of aggregator parameters through training, while the rightmost column shows the obtained heterogeneity gain. This result empirically demonstrates that HED rediscovers the reward structure predicted by our theory to maximize the gain, making the inner aggregator convex, and the outer aggregator concave. We report mean and standard deviation for 90M training frames over 13 random seeds.
@@ 6 Discussion
This work introduces tools for both diagnosing and designing reward functions that incentivize heterogeneity in cooperative MARL. In task allocation settings, our theory shows that the advantage of behavioral diversity is a predictable consequence of reward curvature: if the inner aggregator is Schur-convex, amplifying inequality, and the outer aggregator is Schur-concave, amplifying uniformity, heterogeneous policies are strictly superior; reversing the curvature removes the benefit. Complementing this analysis, and covering settings where our theory doesn’t apply, the proposed HED algorithm automatically steers underspecified environments to either side of the diversity boundary, letting us encourage or suppress heterogeneity and providing a sandbox for studying its advantages. Together, these results help turn the choice of heterogeneity from an ad-hoc heuristic into a controllable design dimension, and help reconcile past mixed results on parameter sharing.
A key remaining open question concerns how the environment’s transition dynamics interact with reward curvature to shape heterogeneity gains. It is also important to consider how one can extend our findings to inseparable tasks. We expand on these open questions, and other limitations, in App. L.
Acknowledgments and Disclosure of Funding
This work is supported by European Research Council (ERC) Project 949940 (gAIa) and ARL DCIST CRA W911NF-17-2-0181. We gratefully acknowledge their support.

### v2

@@ 3 Analysis
4 Heterogeneity Gain Parameter Search (HetGPS)
In complex scenarios where theory might be less applicable, we study heterogeneity through algorithmic search. We consider the setting of a Parametrized Dec-POMDP (PDec-POMDP, defined in App. O). A PDec-POMDP represents a Dec-POMDP (Oliehoek et al., 2016), where the observations, transitions, or reward are conditioned on parameters $\theta$. Hence, the return obtained by the agents, $G^{\theta}(\mathbf{\pi})$, can be differentiated with respect to $\theta$: $\nabla_{\theta}G^{\theta}(\mathbf{\pi})=\frac{\partial}{\partial_{\theta}}G^{\theta}(\mathbf{\pi})$. In particular, computing this gradient in a differentiable simulator allows us to back-propagate through time and optimize $\theta$ via gradient ascent [^3] Although the same approach can train policies (Xu et al., 2022; Song et al., 2024), HetGPS instead optimizes environment parameters and policies separately, using standard zeroth-order policy-gradient methods, to avoid being trapped in local minima. The overhead of implementing HetGPS in this way is modest: it increased training time by roughly 25% in our Sec. 5 experiments compared to training on a fixed environment..
Algorithm 1 Heterogeneity Gain Parameter Search (HetGPS)
@@ 4 Heterogeneity Gain Parameter Search (HetGPS)
4:  HetGainθ = ComputeGain($_{\mathrm{het}}^{\theta},Batch$$_{\mathrm{hom}}^{\theta}$)
5:  if train_env(ii) then
@@
7:  if train_agents(ii) then
@@
Heterogeneity Gain Parameter Search (HetGPS). We now consider the problem of learning the environment parameters $\theta$ to maximize the empirical heterogeneity gain. The empirical heterogeneity gain is defined as the difference in performance between heterogeneous and homogeneous teams in a given PDec-POMDP parametrization. We compare neurally heterogeneous agents (independent parameters) with neurally homogeneous agents (shared parameters). We denote their policies as $\mathbf{\pi}_{\mathrm{het}}$ and $\mathbf{\pi}_{\mathrm{hom}}$. Then, we can simply write the gain as: $\mathrm{HetGain}^{\theta}(\mathbf{\pi}_{\mathrm{het}},\mathbf{\pi}_{\mathrm{hom}})=G^{\theta}(\mathbf{\pi}_{\mathrm{het}})-G^{\theta}(\mathbf{\pi}_{\mathrm{hom}})$, representing the return of heterogeneous agents minus that of homogeneous agents on environment parametrization $\theta$. HetGPS, shown in Alg. 1, learns $\theta$ by performing gradient ascent to maximize the gain: $\theta\leftarrow\theta+\alpha\nabla_{\theta}\mathrm{HetGain}^{\theta}(\mathbf{\pi}_{\mathrm{het}},\mathbf{\pi}_{\mathrm{hom}})$. The environment and the agents are trained in an iterative, bilevel optimization process. We discuss this process, and alternatives when the simulator is non-differentiable, in App. P. At every training iteration, HetGPS collects roll-out batches in the current environment $\theta$ for both heterogeneous and homogeneous teams, computing the heterogeneity gain on the collected data. Then, it updates $\theta$ to maximize the heterogeneity gain. Finally, to train the agents, it uses MARL, with any on-policy algorithm (e.g., MAPPO (Yu et al., 2022)). The functions train_env and train_agents determine when to train each of the components in HetGPS. We consider two possible training regimes: (1) alternated: where HetGPS performs cycles of $x$ agent training iterations followed by $y$ environment training iterations and (2) concurrent: where agents train at every iteration and the environment is updated every $x$ iterations. Note that by performing descent instead of ascent, HetGPS can also be used to minimize the heterogeneity gain.
@@ 5 Experiments
To empirically ground our theoretical analysis, we conduct a three-stage experimental study in cooperative MARL. We first analyze a one-step, observation-free matrix game in which each agent allocates effort $r_{ij}$ over $M$ tasks, and consider reward structures defined by aggregator pairs $U,T\in\{\min,\mathrm{mean},\max\}$. We find that the agents’ learned policies recover the exact heterogeneity gains derived in the theory (Fig. 2). Next, we transfer the same reward structures into embodied, time-extended environments: Multi-goal-capture, 2v2 tag, and football. We show that our curvature theory continues to be and informative in these settings as well. We discuss the learning dynamics that result, and perform further experiments highlighting the difference between neural and behavioral heterogeneity (Bettini et al., 2023), important for understanding our insights. Finally, to study HetGPS, we parametrize the reward structure of Multi-goal-capture using either parametrized Softmax or Power-Sum aggregators (App. I), and run HetGPS to learn parameterizations that maximize the heterogeneity gain. HetGPS learns the theoretically optimal aggregator instantiations, validating its effectiveness at discovering heterogeneous missions. Implementation details and visualizations are available in App. C and App. K.
@@
(i) Task Allocation. We consider a one-step observationless matrix game where $N$ agents need to choose between $M$ tasks. Their actions are effort allocations $r_{ij}$ with $r_{ij}\geq 0$$\sum_{j}r_{ij}=1$, composing matrix $\mathbf{A}$. With aggregators taken from the set $U,T\in\{\min,\mathrm{mean},\max\}$, our goal is to empirically confirm the heterogeneity gains derived in the theory in a learning context. Each time the game is played, all agents obtain the global reward $R(\mathbf{A})$ computed through the double aggregator. We consider two setups: (1) Continuous ($r_{ij}\in\mathbb{R}_{0\leqslant x\leqslant 1}$): agents can distribute their efforts across tasks, (2) Discrete ($r_{ij}\in\{0,1\}$): agents choose only one task. We train with $N=M=4$ for 12 million steps. Fig. 6 shows the evolution of the heterogeneity gains. The final results match exactly the theoretical predictions of Fig. 2 and our curvature theory: concave outer and convex inner aggregators favor heterogeneity. Additional details and results, e.g., for $N=M=2,8$, are in App. J.
@@
Figure 3: Heterogeneity gain for the discrete and continuous matrix games with $N=M=4$ over training iterations. We report mean and standard deviation after 12M frames over 9 random seeds. The final results match the theoretical predictions in the Table of Fig. 2. Solid lines indicate reward structures predicted by theory to have $\Delta R>0$ in either the discrete or continuous setting; dashed lines indicate predicted no gain in both settings.
(ii-1) Multi-goal-capture. Next, we investigate a time-extended, embodied scenario called Multi-goal-capture, based on multi-goal navigation missions (Terry et al., 2021). In Multi-goal-capture, agents need to navigate to goals, and efforts $r^{t}_{ij}$ are continuous scalars computed based on their proximity to these goals. We provide details in App. K. Our goal is to show that the results obtained in the matrix game still hold in this embodied, long-horizong setting. We again consider aggregators $U,T\in\{\min,\mathrm{mean},\max\}$. After 30M training frames (Fig. 4(a)) the empirical heterogeneity gains differ, numerically, from those of the static matrix-game because agents now realize their allocations $r_{ij}$ through time-extended motion. Nonetheless, our curvature theory reliably predicts when there is a heterogeneity gain (Fig. 2): it is positive only for the concave–convex pairs $U=\min,\,T=\max$ and $U=\mathrm{mean},\,T=\max$. We further explain these results (including the interesting presence of “negative” heterogeneity gains) in App. K. Note that the aggregator pairs in this experiment are not contrived: they encode practically meaningful global objectives. For example, $U=\max,T=\max$ implies “at least one agent should go to at least one goal”; $U=\max,T=\min$ implies “all agents should go to the same goal”, and so on. $U=\min,T=\max$, a concave-convex setting shown by our theory to favor heterogeneity, implies “each agent should go to a different goal and all goals should be covered” which is a natural goal for this scenario. This is because $T=\max$ encodes a task that needs just one agent to be completed (e.g., find an object), while $U=\min$ encodes that all tasks should be attended (i.e., agents need to diversify their choices).
(a) Multi-goal-capture.
(b) 2v2 Tag.
@@
Figure 4: Heterogeneity gain for Multi-goal-capture and 2v2 Tag throughout training. We report mean and standard deviation for 30 million training frames over 9 random seeds.
Observability-Heterogeneity Trade-Off: To understand our theoretical results, it is important to solidify the difference between neural heterogeneity (agents having different neural networks) and behavioral heterogeneity (agents acting differently). Our insights concern behavioral heterogeneity, which need not be neurally induced. We show this in App. N, showing that: as the observability of neurally homogeneous agents increases (allowing them to sense each other), these agents can become behaviorally heterogeneous, and thus optimize the heterogeneity gain. This result is visualized here.
(ii-2) 2v2 tag. In our tag experiment, two learning chasers pursue two heuristic escapees in a randomized obstacle field. We define the effort $r_{ij}^{t}$ to be $1$ if chasing agent $i$ manages to capture escaping agent $j$ by time $t$, and 0 otherwise. Whereas Multi-goal-capture had continuous effort allocations, here they are discrete. The global reward is again computed with aggregators $U,T\in\{\min,\text{mean},\max\}$, and is awarded at every time step. This is a sparse reward signal only awarded upon mission success. For example, $U=\min,T=\max$ pays out only if both escapees are each caught by a chaser, encouraging heterogeneity. Training outcomes are summarized in 4(b). We again see that our theoretical results in Fig. 2 (discrete efforts) predict exactly which aggregators will exhibit $\Delta R>0$. More details and visuals are available in App. L and here.
(ii-3) Football. We evaluate our theory in a complex continuous control football game to explore what happens when our reward structure $R(A)$ is only part of a global cooperative reward. To this end, we design a drill in the VMAS Football environment (Bettini et al., 2022), where one agent is tasked to score, while the other has to block the incoming opponent. App. M shows that, also in this case, our theory is highly predictive, with visuals available here.
@@
(iii) Heterogeneous Reward Design. We apply HetGPS to Multi-goal-capture, and ask whether it finds the same aggregator parameterizations predicted by our theory to maximize the heterogeneity gain. We turn the environment into a PDec-POMDP by parameterizing the reward as ${R}^{\theta}(\mathbf{A}^{t})=\bigoplus_{j=1}^{M}\!{}^{\theta}\bigoplus_{i=1}^{N}\!{}^{\theta}r^{t}_{ij}$, with parametrized inner and outer aggregators $U^{\theta}=\bigoplus^{\theta},T^{\theta}=\bigoplus^{\theta}$. Our goal is to learn the parameters $\theta=(\tau_{1},\tau_{2})$, parametrizing $T$ and $U$ respectively, that maximize the heterogeneity gain. We consider two parametrized aggregators from Tab. 3: Softmax and Power-Sum. Softmax: we parameterize both $U^{\theta}$ and $T^{\theta}$ using Softmax. We initialize $\tau_{1}=\tau_{2}=0$, so $U$ and $T$ are initially mean, and run HetGPS (in App. Q we show that HetGPS is robust to adversarial initializations). In 5(a), we show that, to maximize the heterogeneity gain, HetGPS learns to maximize $\tau_{1}$, making $T$ Schur-convex, while minimizing $\tau_{2}$, making $U$ Schur-concave. Hence, it rediscovers the theoretically optimal reward function. The large variance in final parameters occurs because the Softmax aggregator saturates for large magnitudes (e.g., $|\tau|>5$); HetGPS correctly identifies this, leading seeds to converge to arbitrary large values within it. Power-Sum: we parametrize both aggregators with Power-Sum. We initialize both functions to $\tau_{1}=\tau_{2}=1$, representing sum, and run HetGPS. We constrain $\tau_{1,2}\in[0.3,6]$ to stabilize learning. In 5(b), we show that HetGPS learns to maximize $\tau_{1}$, making $T$ Schur-convex, while minimizing $\tau_{2}$, making $U$ Schur-concave; again rediscovering the optimal parametrization our theory predicts. These results simultaneously validate HetGPS and our curvature theory, since each arrives at the same reward structure independently.
@@
(a) HetGPS with Softmax aggregators ($\tau\in\mathbb{R}$). $\tau\geq 0$ is Schur-convex; $\tau\leq 0$ is Schur-concave.
(b) HetGPS with Power-Sum aggregators ($\tau\in[0.3,6]$). $\tau\geq 1$ is Schur-convex; $\tau\leq 1$ is Schur-concave.
@@
Figure 5: HetGPS results in Multi-goal-capture. The two leftmost columns report the evolution of aggregator parameters through training, while the rightmost column shows the obtained heterogeneity gain. This result empirically demonstrates that HetGPS rediscovers the reward structure predicted by our theory to maximize the gain, making the inner aggregator convex, and the outer aggregator concave. We report mean and standard deviation for 90M training frames over 13 random seeds.
@@ 6 Discussion
This work introduces tools for both diagnosing and designing reward functions that incentivize heterogeneity in cooperative MARL. In task allocation settings, our theory shows that the advantage of behavioral diversity is a predictable consequence of reward curvature: if the inner aggregator is Schur-convex, amplifying inequality, and the outer aggregator is Schur-concave, amplifying uniformity, heterogeneous policies are strictly superior; reversing the curvature removes the benefit. Complementing this analysis, and covering settings where our theory doesn’t apply, the proposed HetGPS algorithm automatically steers underspecified environments to either side of the diversity boundary, letting us encourage or suppress heterogeneity and providing a sandbox for studying its advantages. Together, these results help turn the choice of heterogeneity from an ad-hoc heuristic into a controllable design dimension, and help reconcile past mixed results on parameter sharing.
A key remaining open question concerns how the environment’s transition dynamics interact with reward curvature to shape heterogeneity gains. We expand on this open question, other directions for future work, and our scope/limitations, in App. R.